OpenStudy (anonymous):

Let f(x) = 3 x^2 - 8. Find an equation of the line tangent to the graph of f at the point where x = -3.

OpenStudy (anonymous):

what is the equation first?

OpenStudy (anonymous):

first find the second coordinate by replacing x by -3

OpenStudy (mathhelp346):

the equation would be 3 -3^2-8

OpenStudy (anonymous):

y=14

OpenStudy (anonymous):

i get \[y=3\times 9-8=27-8=19\] so the point on your line is \[(-3,19)\] then find the slope. the derivative is \[y'=6x\] and replacing x by -3 gives you \[-18\] so you have a slope and a point, use point slope formula

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

\[y-19=-18(x+3)\] etc

OpenStudy (mertsj):

Well imagine that. The same answer as when I helped you with the same problem.

OpenStudy (anonymous):

yeah but its not coming out right

OpenStudy (mertsj):

What do you mean it's not coming out right?

OpenStudy (anonymous):

I plugged in y=-18x+b and y-19=-18(x+3) and on the website Im supposed to submit is saying that it is incorrect

OpenStudy (mertsj):

Well plug in y = -18x-35

OpenStudy (mertsj):

You told me before that your book said the answer was y = -18x+b

OpenStudy (anonymous):

I know but the website wasnt taking that answer

OpenStudy (mertsj):

Did you put in y = -18x-35 yet?

OpenStudy (anonymous):

but the website just took that answer so its right and the book didnt want me to simplify

OpenStudy (anonymous):

thanx

OpenStudy (mertsj):

Well i think I told you that was the answer 1 1/2 hours ago.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

what a waste of 1 1/2 hours